OpenStudy (anonymous):

6/y+1= 5/y-7 /olve for y??

jhonyy9 (jhonyy9):

y+1 and y-7 are denominators all ?

jhonyy9 (jhonyy9):

if just y is denominator you will get 6+y=5-7y 8y=-1 y= -1/8